A hydrocarbon has a weight of 78 gives carbon dioxide and water on combustion with 3.385 of carbon and 0.682 of water find the formula


1
Expert's answer
2020-09-29T06:16:26-0400

Calculate the moles of C and of H formed in the combustion:


"\dfrac{3.385}{44}=0.0769 moles" CO2 * ("\dfrac{1 mole}{1moles_{CO_{2}}})" "=0.0769moles" of C


"\dfrac{0.682}{18}=0.0379moles" H2O * ("\dfrac{2 mole}{1moles_{H_{2}O}})" = "0.0758moles" of H


Divide each by the smaller:


C = "\dfrac{0.0769}{0.0758}=1.0"


H = "\dfrac{0.0758}{0.0758} = 1.0"


The empirical formula is CH

Since the empirical formula mass is 13 "( \tfrac{g}{mole})" dividing the molar mass of the compound by the formula mass gives "\\tfrac{78}{16}\t=6" .


The molecular formula is C6H6